You draw a map with a scale of 10 miles to 0.5 inch. How far should you space two cities that are 60 miles apart?
Answer by jim_thompson5910(35256)

(10 miles)/(0.5 inch) = (60 miles)/(x inches)


10/0.5 = 60/x


10x = 0.5*60


10x = 30


x = 30/10


x = 3


You should space them 3 inches apart.